OpenStudy (anonymous):

A regular polygon has 7 sides. the measure of each interior angle is_____? and the measure of each exterior angle is_____? (7-2) 180=900/7

OpenStudy (anonymous):

900/7 = each interior angle for exterior angle, just take 180 - interior angle

OpenStudy (anonymous):

Interior: In radians, (1-2/n) x pi, n = 7 In degrees, 180 - 360/n degrees, n = 7

OpenStudy (anonymous):

You can also think of the exterior angles all adding up to 360 degrees (or 2 pi) as you go around the polygon, you start and finish in the same direction. Therefore you can also figure out the exterior angle by dividing 2pi/n or 360/n for radians or degrees respectively.

OpenStudy (anonymous):

The extra geeky answer ;o)

OpenStudy (anonymous):

 find x and y